I need help solving this equation that I've been tryin to solve for a while now. can someone please help?
If you were to use the Gauss-Jordan elimination on the following linear system, what would the augmented matix be?
3w-x=2y+z-4
9x-y+z=10
4w+3y-z=7
12x+17=2y-z+6?

Answer by Mathtut(3670)

lets re write the equations and line up the variables from each equation.
:
3w-1x-2y-z=-4
0w+9x-1y+z=10
4w+0x+3y-z=7
0w+12x-2y+z=-11
: